# Autocovariance and Optimal Design for Random Walk Metropolis-Hastings Algorithm

### v1

@@ Abstract
The Metropolis–Hastings algorithm has been extensively studied in the estimation and simulation literature, with most prior work focusing on convergence behavior and asymptotic theory. However, its covariance structure—an important statistical property for both theory and implementation—remains less understood. In this work, we provide new theoretical insights into the scalar case, focusing primarily on symmetric unimodal target distributions with symmetric random walk proposals, where we also establish an optimal proposal design. In addition, we derive some more general results beyond this setting. For the high-dimensional case, we relate the covariance matrix to the classical 0.23 average acceptance rate tuning criterion.
@@ I Introduction
Markov chain Monte Carlo (MCMC) provides a general framework for sampling from complex distributions by constructing a Markov chain with the desired stationary distribution [Spall2003]. This approach enables efficient approximation of expectations and integrals that are otherwise analytically intractable, and it has found widespread applications in statistics, engineering, and control [HillandSpall]. In particular, Monte Carlo particle filters can discretize the state space to enable efficient MAP sequence estimation via dynamic programming [Godsill2001]. MCMC has also been applied to Bayesian inference for stochastic kinetic models by sampling joint posteriors of parameters and latent data [Golightly2005], and to system identification for Bayesian parameter estimation in complex dynamical models beyond the reach of traditional techniques [Ninness2010]. In the control domain, particle MCMC has been applied to infer latent trajectories and nonparametric dynamics in Gaussian process state-space models [Frigola2013], MCMC has been used to generate model uncertainty sets in nonlinear dynamical systems via Hamiltonian Monte Carlo sampling of high-dimensional transfer operators [Srinivasan], and ghost sampling, an MCMC-based technique, has been introduced to efficiently simulate power disturbances conditional on rate-of-change-of-frequency violations in power systems [Moriarty]; moreover, MCMC has also been integrated with Bernstein approximation techniques to handle non-affine and dependent chance constrained optimization problems, providing an efficient approach to high-dimensional integration [Zhao].
The Metropolis–Hastings algorithm is one of the foundational methods in MCMC, with widespread applications in model estimation, physics, and system identification [HillandSpall, Metropolis1953]. The theoretical study of MH has been extensive, with a large body of work devoted to convergence behavior and asymptotic properties of estimators [Tierney1994, RobertsandRosenthal, BrooksandGareth1998]. An important line of research in MCMC has focused on the scaling and efficiency of Metropolis-Hastings algorithms in high-dimensional settings. For Gaussian random walk Metropolis (RWM) algorithms, the efficiency of the chain is maximized when the average acceptance rate is tuned to approximately $0.23$ [RobertsandGelman]. This “$0.23$ rule” has since become a widely used guideline for practitioners, providing a concrete and interpretable criterion for tuning proposal variances in high-dimensional problems.
@@
This work aims to address this gap. We first study the scalar case of MH with symmetric random walk proposals, focusing on symmetric unimodal target distributions. In this setting, we establish new theoretical results on the covariance structure and propose an optimal design for the proposal distribution of symmetric random walk form. In addition, we present some general results beyond this restrictive setting. For the high-dimensional case, under the same setup as Roberts and Gelman [RobertsandGelman], we relate the unit-lag covariance matrix to the classical 0.23 average acceptance rate criterion, providing new theoretical connections between covariance properties and established tuning principles.
@@ II Preliminaries
Monte Carlo methods provide a general tool for estimation and simulation when analytical solutions are unavailable. MCMC extends this idea by constructing a Markov chain with the desired stationary distribution, enabling sampling from complex or high-dimensional models [Spall2003, HillandSpall]. Within this framework, the Metropolis-Hastings (MH) algorithm stands out as a versatile and widely used scheme. In this section, we present the Metropolis-Hastings algorithm, establish the notation, and outline the problem setup.
@@ II-A Metropolis-Hastings algorithm
where $\delta(\cdot)$ is the Dirac measure at 0, and $r(x)=\int\alpha(x,y)\,q(y|x)\,\mathrm{d}y$ is the overall acceptance probability at current state [HillandSpall]. Under ergodicity and irreducibility assumptions, the Markov chain converges to the unique invariant distribution. Let $(\mathcal{X},\mathcal{B})$ denote the underlying measurable state space, with $\mathcal{X}$ the state space and $\mathcal{B}$ its associated $\sigma$-algebra, then
@@ II-B Notation
Throughout this paper we assume that the target distribution $\pi$ admits a density with respect to Lebesgue measure. Similarly, the proposal distribution $q(\cdot|x)$ is assumed to admit a density $q(y|x)$ with respect to the corresponding base measure. $\Pi(\cdot)$ is the cumulative distribution function corresponding to $\pi(\cdot)$, and we will use the cdf notation for brevity in what follows. Unless otherwise stated, all integrals are taken over $\mathbb{R}$ with respect to Lebesgue measure. We also adopt the standard stochastic order notation: $Z_{n}=o_{p}(a_{n})$ means $Z_{n}/a_{n}\to 0$ in probability, while $Z_{n}=O_{p}(a_{n})$ means $Z_{n}/a_{n}$ is bounded in probability. Also for simplicity, let $\partial_{x}f(x)$ and $\partial_{x}^{2}f(x)$ be the first and second derivatives of a function $f(x)$ with respect to its argument. Lastly, for a matrix $A$, $\|A\|_{\mathrm{op}}\mathrel{\mathop{\ordinarycolon}}=\max\{\,\|Ax\|_{2}\mathrel{\mathop{\ordinarycolon}}x\in\mathbb{R}^{n},\ \|x\|_{2}=1\,\}$ denotes its operator (spectral) norm.
@@ II-C Problem Setup
Consider the unit-lag covariance of the output sequence of the MH algorithm under the stationarity assumption, i.e., the current state $X_{t}\sim\pi(\cdot)$, which implies that the next state $X_{t+1}\sim\pi(\cdot)$ as well. Stationarity holds provided that the chain is irreducible, aperiodic, and satisfies detailed balance with respect to $\pi(\cdot)$, ensuring that $\pi(\cdot)$ is the unique invariant distribution [HillandSpall]; these conditions are known to hold for an MH algorithm.
@@ III Main results
Then by reversibility, i.e. $\pi(x)q(y|x)=\pi(y)q(x|y)$,
@@
Applying the Tonelli’s theorem [Tonelli] to justify the interchange of integrals in (III) since the integrand $(x-y)^{2}\alpha(x,y)q(y|x)\pi(x)$ is nonnegative, and using stationarity, we obtain (3.1). ∎
@@
Recalling that the proposal $q(\cdot|x)$ is of the random walk form, $q(y|x)=\phi(y-x)$ and $q(x|y)=\phi(x-y)$. Thus by definition, the acceptance rate under the shifted target density $\pi_{c}(\cdot)$ is
@@
Let $\phi(x)$ be any probability density that is symmetric about 0. Then,
$|x|\phi(x)<\frac{1}{2}\text{ a.s.}$ (5)
@@
Assume $\exists\,x_{0}\in\mathbb{R}$ s.t.
$|x_{0}|\phi(x_{0})\geq\frac{1}{2}\Rightarrow\phi(x_{0})\geq\frac{1}{2|x_{0}|}.$
By symmetry, $\phi(-x_{0})=\phi(x_{0})$, so
$\int_{-|x_{0}|}^{|x_{0}|}\phi(x)\,dx\geq\int_{-|x_{0}|}^{|x_{0}|}\frac{1}{2|x_{0}|}\,dx=1.$
Since $\phi$ is a probability density, equality must hold and $\phi(x)=\frac{1}{2|x_{0}|}$ a.s. on $|x_{0}|,|x_{0}|]$, i.e., $\phi$ is uniform, that is,
$|x|\phi(x)=\frac{|x|}{2|x_{0}|}<\frac{1}{2},\quad\forall\,|x|<|x_{0}|.$
At $x=\pm|x_{0}|$, equality may hold pointwise, but probability densities are defined up to sets of measure zero. We reached a contradiction. ∎
Lemma 3.3 is powerful in that it holds for any symmetric probability density, including standard distributions such as the normal $\mathcal{N}(0,1)$, for which $\max_{x}\{|x|\phi(x)\}=e^{-\frac{1}{2}}/{\sqrt{2\pi}}<1/2$. Its general validity makes it an essential tool for the proof of our main theorem.
@@
where the second equality follows from a change of variables together with the Tonelli’s theorem [Tonelli] and the third equality follows from the fact that $f(\cdot)$ is even.
Given the target density is symmetric unimodal, now consider $f(x),\,x\geq 0$
@@
By Corollary 3.2, proposal densities of symmetric random–walk form $q(y|x)=\phi(|y-x|)$ is transition-invariant. Then the covariance admits the same form as in the mean–zero case, with $\Pi$ replaced by $\Pi_{\mu}$. ∎
@@ IV Optimal Proposal Design
Note that $\arg\max$ may define a set, though for symmetric unimodal target densities it commonly reduces to a unique maximizer.
Now we check if such $y^{*}$ exists. First, the finite variance assumption implies $\mathbb{E}[X^{2}]<\infty$, which ensures
@@
Indeed, since $1-\Pi(x)=\mathbb{P}(X>x)\leq\mathbb{E}[X^{2}]/x^{2}$ by Markov’s inequality, the limit can only be 0, a strictly positive constant, or $\infty$. The latter two would force $\mathbb{E}[X^{2}]=\infty$, contradicting the finite variance assumption.
From (11), and noting that $w(0)=0$, $w(y)>0$ for all $y>0$, $\lim_{y\to\infty}w(y)=0$, and $w(\cdot)$ is continuous, the Extreme Value Theorem guarantees that both the maximum and the minimum (zero) exist. Hence $w_{\max}\mathrel{\mathop{\ordinarycolon}}=w(y^{*})$ must exist for any non-degenerate target distribution with finite variance.
@@
with the additional requirement that the second derivative at $y^{*}$ is negative to ensure optimality.
@@
equation (12) has a unique solution when the target density $\pi(\cdot)$ is log-concave. In this case, define $g(y)\mathrel{\mathop{\ordinarycolon}}=\pi(y)/(1-\Pi(y))$, which is monotone increasing, implying that $h(y)=yg(y)$ is strictly increasing. Hence the equation $h(y)=2$ has a unique solution, and since $w(0)=w(\infty)=0$. Moreover, this critical point is necessarily the global maximizer of $w(y)$. Indeed, if it were a saddle point, then since $w(x)$ is continuously differentiable, additional critical points would have to exist, which contradicts the uniqueness.
However, under the regularity condition, the optimization problem becomes an infimum problem rather than a maximization problem. In this case, the maximum achieved by the two-point measure cannot be attained, since the regularity condition is violated. We can prove this rigorously, as shown in Proposition 4.2.
@@
By symmetry, with $h(y)\mathrel{\mathop{\ordinarycolon}}=2\phi(2y)$ we have
@@
with equality iff $h$ is supported on $\mathcal{M}$. Thus any symmetric atomic measure supported on $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}y\in\mathcal{M}\}$ achieves $w_{\max}$.
If $\phi$ is a probability density, then the induced function $h(y)=2\phi(2y)$ is continuous in the sense that it cannot concentrate all its probability mass at finitely many points. The set $\mathcal{M}$ consists only of isolated points. Then away from $\mathcal{M}$, the continuity of $w(\cdot)$ guarantees that there exists $\varepsilon>0$ such that $w_{\max}-w(y)\geq\varepsilon$ for all $y$ sufficiently far from $\mathcal{M}$. Since any density $h$ must assign positive probability to regions outside arbitrarily small neighborhoods of $\mathcal{M}$, the weighted average $\int_{0}^{\infty}w(s)h(s)\,\mathrm{d}s\leq w_{\max}/2-\varepsilon/2<w_{\max}/2$, and hence $\mathcal{J}(\phi)<w_{\max}$.
Nevertheless, the two-point measure can be approximated arbitrarily well by densities concentrated near the $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}y\in\mathcal{M}\}$, such as narrow Gaussians. As the concentration increases, the corresponding values of $\mathcal{J}(\phi)$ approach $w_{\max}$. Hence $\sup\mathcal{J}=w_{\max}$, but the maximum is not attained within the class of densities when the maximizers are isolated points.
@@
Continuing with the analysis of the unit-lag covariance structure, we next relate it to the “0.23 rule” for random-walk Gaussian Metropolis algorithms via the unit-lag covariance matrix. This classic criterion arises from diffusion limit analysis in high dimensions [RobertsandGelman] and prescribes tuning the average acceptance probability to about $0.23$ for optimal efficiency. Here we reinterpret this criterion through the lens of the unit-lag covariance matrix. We now state Theorem 4.3, following the notation and the setup of Gelman and Roberts [RobertsandGelman].
@@
where each $p_{i}$ is a strictly positive $C^{2}$ density on $\mathbb{R}$ with $\sigma_{i}^{2}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Var}_{p_{i}}(X_{i})\in(0,\infty),\,m_{i}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{p_{i}}\big[(\partial_{x}\log p_{i}(X_{i}))^{2}\big]\in(0,\infty),$ and $\mathrm{E}_{p_{i}}[|\partial_{x}\log p_{i}(X_{i})|^{3}],\,\mathrm{E}_{p_{i}}[|\partial_{x}^{2}\log p_{i}(X_{i})|]<\infty$ for all $i$. Consider a stationary random-walk Metropolis chain with the symmetric random walk Gaussian proposal density, i.e. $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{z}$, where $\boldsymbol{z}\sim\mathcal{N}\!\Big(\boldsymbol{0},\tfrac{\ell^{2}}{d}\,\boldsymbol{I}_{d}\Big)$, $\forall\ell>0$, and $\boldsymbol{I}_{d}$ denotes the $d\times d$ identity matrix. Define the averages $\bar{\sigma}_{d}^{2}\mathrel{\mathop{\ordinarycolon}}=d^{-1}\sum\limits_{i=1}^{d}\sigma_{i}^{2},\,\bar{m}_{d}\mathrel{\mathop{\ordinarycolon}}=d^{-1}\sum\limits_{i=1}^{d}m_{i}.$ Denote by $\bar{m}\mathrel{\mathop{\ordinarycolon}}=\lim\limits_{d\to\infty}\bar{m}_{d}$, assuming the limit exists in $(0,\infty)$. Then the unit-lag covariance matrix satisfies
@@
By Gelman and Roberts and apply the change of variable, let $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{z}$,
@@
Then fix $\boldsymbol{x}$ and set $f(\boldsymbol{x,z})\mathrel{\mathop{\ordinarycolon}}=\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})$. We look at $\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\boldsymbol{z}\boldsymbol{z}^{\top}]$, for $\boldsymbol{z}\sim\mathcal{N}(\boldsymbol{0},\ell^{2}\boldsymbol{I}_{d}/d)$, Applying Stein’s Lemma [stein] and the assumption $\mathrm{E}_{p_{i}}[|\partial_{x}^{2}\log p_{i}(X_{i})|]<\infty$ gives,
@@
$\tfrac{1}{2}\Big(\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})\,\boldsymbol{z}\boldsymbol{z}^{\top}]-\tfrac{\ell^{2}}{d}\,\mathbb{E}[\alpha(\boldsymbol{x},\boldsymbol{x}+\boldsymbol{z})]\boldsymbol{I}_{d}\Big),$
it follows that $\boldsymbol{R}_{d}$ is real symmetric, hence Hermitian as well. Therefore both the main diagonal term and the perturbation matrix $\boldsymbol{R}_{d}$ are Hermitian. Then by Weyl’s inequality [Horn_Johnson_2012], for Hermitian matrices $\boldsymbol{A}$ and $\boldsymbol{E}$,
@@
$\lambda_{\min}\!\big(\mathrm{Cov}(\boldsymbol{x}_{t},\boldsymbol{x}_{t+1})\big)\;\geq\;\min_{i}\!\Big\{\sigma_{i}^{2}-\tfrac{\ell^{2}}{d}\cdot\Phi(-\tfrac{1}{2}\ell\sqrt{\bar{m}})\Big\}-\|\boldsymbol{R}_{d}\|_{\mathrm{op}}.$
Since $\min\limits_{i}\sigma_{i}^{2}>0$, the subtracted terms are of order $O(d^{-1})$ and $o(d^{-1})$, hence strictly smaller than $\min\limits_{i}\sigma_{i}^{2}$ for sufficiently large $d$. This shows that the covariance matrix is positive definite in high dimension. ∎
Theorem 4.3 shows that in the high-dimensional regime, for a product-form target with non-identically distributed components, the diagonal entries of the unit-lag covariance admit the expansion as equation (19). Gelman and Roberts define the follwoing

### v2

# arXiv:2511.10967v2[stat.CO] 28 Feb 2026

@@ Abstract
The Metropolis–Hastings algorithm has been extensively studied in the estimation and simulation literature, with most prior work focusing on convergence rates and asymptotic theory. In contrast, the covariance structure of the resulting Markov chain—an important determinant of statistical efficiency—has received comparatively little theoretical attention. In this work, we study the covariance properties of Metropolis–Hastings algorithms in both scalar and high-dimensional settings. For the scalar case, we focus on symmetric unimodal target distributions with symmetric random-walk proposals, and prove positivity of the unit-lag covariance, together with explicit structural characterizations. Within this class, we further establish an optimal proposal design that minimizes covariance and improves sampling efficiency. In the high-dimensional regime, we relate the covariance matrix to the classical 0.23 average acceptance-rate tuning criterion, providing a new covariance-based perspective on optimal scaling.
@@ I Introduction
Markov chain Monte Carlo (MCMC) provides a general framework for sampling from complex distributions by constructing a Markov chain with the desired stationary distribution [15]. This approach enables efficient approximation of expectations and integrals that are otherwise analytically intractable, and it has found widespread applications in statistics, engineering, and control [6]. In particular, Monte Carlo particle filters can discretize the state space to enable efficient MAP sequence estimation via dynamic programming [4]. MCMC has also been applied to Bayesian inference for stochastic kinetic models by sampling joint posteriors of parameters and latent data [5], and to system identification for Bayesian parameter estimation in complex dynamical models beyond the reach of traditional techniques [12]. In the control domain, particle MCMC has been applied to infer latent trajectories and nonparametric dynamics in Gaussian process state-space models [3], MCMC has been used to generate model uncertainty sets in nonlinear dynamical systems via Hamiltonian Monte Carlo sampling of high-dimensional transfer operators [16], and ghost sampling, an MCMC-based technique, has been introduced to efficiently simulate power disturbances conditional on rate-of-change-of-frequency violations in power systems [10]; moreover, MCMC has also been integrated with Bernstein approximation techniques to handle non-affine and dependent chance constrained optimization problems, providing an efficient approach to high-dimensional integration [18].
The Metropolis-–Hastings algorithm is one of the foundational methods in MCMC, with widespread applications in model estimation, physics, and system identification [6, 2]. The theoretical study of MH has been extensive, with a large body of work devoted to convergence behavior and asymptotic properties of estimators [17, 14, 1]. An important line of research in MCMC has focused on the scaling and efficiency of Metropolis-Hastings algorithms in high-dimensional settings. For Gaussian random walk Metropolis–Hastings (RWMH) algorithms, the efficiency of the chain is maximized when the average acceptance rate is tuned to approximately $0.23$ [13]. This “$0.23$ rule” has since become a widely used guideline for practitioners, providing a concrete and interpretable criterion for tuning proposal variances in high-dimensional problems.
@@
This work aims to address this gap. We first study the scalar case of MH with symmetric random walk proposals, focusing on symmetric unimodal target distributions. In this setting, we establish new theoretical results on the covariance structure and propose an optimal design for the proposal distribution of symmetric random walk form. In addition, we present some general results beyond this restrictive setting. For the high-dimensional case, under the same setup as Roberts, Gelman and Gilks [13], we relate the unit-lag covariance matrix to the classical 0.23 average acceptance rate criterion, providing new theoretical connections between covariance properties and established tuning principles.
@@ II Preliminaries
Monte Carlo methods provide a general tool for estimation and simulation when analytical solutions are unavailable. MCMC extends this idea by constructing a Markov chain with the desired stationary distribution, enabling sampling from complex or high-dimensional models [15, 6]. Within this framework, the Metropolis-Hastings (MH) algorithm stands out as a versatile and widely used scheme. In this section, we present the Metropolis-Hastings algorithm, establish the notation, and outline the problem setup.
@@ II-A Metropolis-Hastings algorithm
where $\delta(\cdot)$ is the Dirac measure at 0, and $r(x)=\int\alpha(x,y)\,q(y|x)\,\mathrm{d}y$ is the overall acceptance probability at current state [6]. Under ergodicity and irreducibility assumptions, the Markov chain converges to the unique invariant distribution. Let $(\mathcal{X},\mathcal{B})$ denote the underlying measurable state space, with $\mathcal{X}$ the state space and $\mathcal{B}$ its associated $\sigma$-algebra, then
@@ II-B Notation
Throughout this paper we assume that the target distribution $\pi$ admits a density with respect to Lebesgue measure. Similarly, the proposal distribution $q(\cdot|x)$ is assumed to admit a density $q(y|x)$ with respect to the corresponding base measure. $\Pi(\cdot)$ is the cumulative distribution function corresponding to $\pi(\cdot)$, and we will use the cumulative distribution function (CDF) notation for brevity in what follows. Unless otherwise stated, all integrals are taken over $\mathbb{R}$ with respect to Lebesgue measure. We also adopt the standard stochastic order notation: $Z_{n}=o_{p}(a_{n})$ means $Z_{n}/a_{n}\to 0$ in probability, while $Z_{n}=O_{p}(a_{n})$ means $Z_{n}/a_{n}$ is bounded in probability. Also for simplicity, let $\partial_{x}f(x)$ and $\partial_{x}^{2}f(x)$ be the first and second derivatives of a function $f(x)$ with respect to its argument. Lastly, for a matrix $A$, $\|A\|_{\mathrm{op}}\mathrel{\mathop{\ordinarycolon}}=\max\{\,\|Ax\|_{2}\mathrel{\mathop{\ordinarycolon}}x\in\mathbb{R}^{n},\ \|x\|_{2}=1\,\}$ denotes its operator (spectral) norm.
@@ II-C Problem Setup
Consider the unit-lag covariance of the output sequence of the MH algorithm under the stationarity assumption, i.e., the current state $X_{t}\sim\pi(\cdot)$, which implies that the next state $X_{t+1}\sim\pi(\cdot)$ as well. Stationarity holds provided that the chain is irreducible, aperiodic, and satisfies detailed balance with respect to $\pi(\cdot)$, ensuring that $\pi(\cdot)$ is the unique invariant distribution [6]; these conditions are known to hold for an MH algorithm.
@@ III Main results
Then by reversibility, i.e. $\pi(x)q(y|x)\alpha(x,y)=\pi(y)q(x|y)\alpha(y,x)$,
@@
Applying the Tonelli’s theorem [11] to justify the interchange of integrals in (III) since the integrand $(x-y)^{2}\alpha(x,y)q(y|x)\pi(x)$ is nonnegative, and using stationarity, we obtain (3.1). ∎
@@
Note that the variance is invariant under translation, i.e., $\sigma^{2}_{\pi_{c}}=\sigma^{2}_{\pi}$. And recalling that the proposal $q(\cdot|x)$ is of the random walk form, $q(y|x)=\phi(y-x)$ and $q(x|y)=\phi(x-y)$. Thus by definition, the acceptance rate under the shifted target density $\pi_{c}(\cdot)$ is
@@
Let $\phi(x)$ be any probability density on $\mathbb{R}$ that is symmetric and unimodal about 0. Then,
$|x|\phi(x)<\frac{1}{2}\quad\text{a.e.}$
@@
Suppose that the inequality $|x|\phi(x)<1/2$ fails on a set of positive measure. This implies that the set is non-empty, so there exists at least one point $x_{0}\neq 0$ such that
$|x_{0}|\phi(x_{0})\geq\frac{1}{2}.$
By symmetry and unimodality, $\phi(|x|)$ is non-increasing for $x\in\mathbb{R}$. Consequently, for any $x$ in the interval $(-|x_{0}|,|x_{0}|)$, we must have $\phi(x)\geq\phi(x_{0})$. Integrating $\phi$ over this interval yields
$\int_{-|x_{0}|}^{|x_{0}|}\phi(x)\,dx\geq\int_{-|x_{0}|}^{|x_{0}|}\phi(x_{0})\,dx=2|x_{0}|\phi(x_{0})\geq 1.$
Since $\phi$ is a probability density function, we know that $\int_{\mathbb{R}}\phi(x)\,dx=1$. The inequality above implies that all the probability mass must be concentrated within $|x_{0}|,|x_{0}|]$, and furthermore, that $\phi(x)$ must be constant on this interval to avoid the integral strictly exceeding 1.
Thus, $\phi$ must be the uniform distribution on $|x_{0}|,|x_{0}|]$. The condition $|x|\phi(x)\geq 1/2$ holds only at the boundaries $x=\pm x_{0}$, which constitute a set of measure zero. This contradicts the original assumption. ∎
Lemma 3.3 is powerful in that it holds for any symmetric unimodal probability density, including standard distributions such as the normal $\mathcal{N}(0,1)$, for which $\max_{x}\{|x|\phi(x)\}=e^{-\frac{1}{2}}/{\sqrt{2\pi}}<1/2$. Its general validity makes it an essential tool for the proof of our main theorem.
@@
where the second equality follows from a change of variables together with Tonelli’s theorem [11] and the third equality follows from the fact that $f(\cdot)$ is even.
Given that the target density is symmetric unimodal, now consider $f(x),\,x\geq 0$
@@
By Corollary 3.2, proposal densities of symmetric random–walk form $q(y|x)=\phi(|y-x|)$ are transition-invariant. Then the covariance admits the same form as in the mean–zero case, with $\Pi$ replaced by $\Pi_{\mu}$. ∎
@@ IV Optimal Proposal Design
Note that $\arg\max$ may define a set, though for symmetric unimodal target densities it commonly reduces to a unique maximizer. Define
$w(y)\mathrel{\mathop{\ordinarycolon}}=y^{2}[1-\Pi(y)].$
We can show the following
@@
Proof of (10).
We have by Markov’s inequality that
$2x^{2}[1-\Pi(x)]=x^{2}\cdot\mathbb{P}(|X|>x)\leq\mathbb{E}[X^{2};|X|>x].$
it suffices to show that $\mathbb{E}[X^{2};|X|>x]\to 0$ as $x\to\infty$. Note that $X^{2}\mathbf{1}_{\{|X|>x\}}\to 0$ almost surely and $0\leq X^{2}\mathbf{1}_{\{|X|>x\}}\leq X^{2},$ with $\mathbb{E}[X^{2}]<\infty$. Hence, by the Dominated Convergence Theorem, $\mathbb{E}[X^{2};|X|>x]\to 0$
From (10), and noting that $w(0)=0$, $\lim_{y\to\infty}w(y)=0$, and $w(\cdot)$ is continuous, the Extreme Value Theorem guarantees that both the maximum and the minimum (zero) exist. Hence $w_{\max}\mathrel{\mathop{\ordinarycolon}}=w(y^{*})$ must exist for any non-degenerate target distribution with finite variance.
@@
with the additional requirement that the second derivative at $y^{*}$ is negative to ensure local optimality.
@@
By Corollary 3.2, we may assume $\mu=0$ without loss of generality. Equation (11) has a unique solution when the target density $\pi(\cdot)$ is log-concave. In this case, define $g(y)\mathrel{\mathop{\ordinarycolon}}=\pi(y)/(1-\Pi(y))$, which is monotonically increasing, implying that $h(y)=yg(y)$ is strictly increasing. Also note that $\lim\limits_{y\to+\infty}h(y)\geq\lim\limits_{y\to+\infty}y\cdot g(0)=\infty$ Hence, the equation $h(y)=2$ admits a solution, and that solution is unique.
However, under the regularity condition, the optimization problem becomes a supremum problem rather than a maximization problem. In this case, the local maximum achieved by the two-point measure cannot be attained, since attaining it would violate the regularity condition. Proposition 4.2 makes this precise by showing that, when $\phi$ is required to be a probability density satisfying the support condition needed for irreducibility, the supremum cannot be attained, although it can be approximated arbitrarily closely.
@@
By symmetry, with $h(y)\mathrel{\mathop{\ordinarycolon}}=2\phi(2y)$ (noting that $\int_{0}^{\infty}h(s)\,\mathrm{d}s=1/2$), we have
@@
with equality iff $h(s)=0$ for almost every $s\notin\mathcal{M}$. Thus any symmetric atomic measure fully supported on $\{\pm 2y\mathrel{\mathop{\ordinarycolon}}y\in\mathcal{M}\}$ achieves $w_{\max}$.
Now consider the case where $\phi$ is a probability density. Then $h$ is absolutely continuous with respect to the Lebesgue measure. Since $\mathcal{M}$ consists only of isolated points, it has Lebesgue measure zero. Consequently, any density $h$ must assign positive mass to the set $\mathcal{M}^{c}=\{y\geq 0\mathrel{\mathop{\ordinarycolon}}y\notin\mathcal{M}\}$. On this set, $w(y)<w_{\max}$ strictly. Therefore,
$\int_{0}^{\infty}w(s)h(s)\,\mathrm{d}s$ $=\int_{\mathcal{M}^{c}}w(s)h(s)\,\mathrm{d}s$
$<w_{\max}\int_{\mathcal{M}^{c}}h(s)\,\mathrm{d}s=\frac{w_{\max}}{2}$
which implies $\mathcal{J}(\phi)<w_{\max}$. Thus, the supremum is not attained within the class of densities.
As the concentration increases, the mass shifts towards $\mathcal{M}$, and $\mathcal{J}(\phi)$ approaches $w_{\max}$. Hence $\sup\mathcal{J}=w_{\max}$. ∎
@@
Continuing with the analysis of the unit-lag covariance structure, we next relate it to the “0.23 rule” for random-walk Gaussian Metropolis algorithms via the unit-lag covariance matrix. This classic criterion arises from diffusion limit analysis in high dimensions [13] and prescribes tuning the average acceptance probability to about $0.23$ for optimal efficiency. Here we reinterpret this criterion through the lens of the unit-lag covariance matrix. We now state Theorem 4.3, following the notation and the setup of Roberts et al. [13].
@@
where each $p_{i}$ is a strictly positive $C^{2}$ density on $\mathbb{R}$ with $\sigma_{i}^{2}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Var}_{p_{i}}(X_{i})\in(0,\infty)$, $m_{i}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{p_{i}}\big[((\log p_{i})^{\prime}(X_{i}))^{2}\big]\in(0,\infty)$, and $\mathbb{E}_{p_{i}}\big[|(\log p_{i})^{\prime}(X_{i})|^{3}\big]<\infty$, $\mathbb{E}_{p_{i}}\big[|(\log p_{i})^{\prime\prime}(X_{i})|\big]<\infty$ for all $i$. Consider a stationary random-walk Metropolis chain with the symmetric random walk Gaussian proposal density, i.e. $\boldsymbol{Y}=\boldsymbol{X}+\boldsymbol{Z}$, where $\boldsymbol{Z}\sim\mathcal{N}\!\big(\boldsymbol{0},\tfrac{\ell^{2}}{d}\,\boldsymbol{I}_{d}\big)$, $\text{for some }0<\ell<\infty$, and $\boldsymbol{I}_{d}$ denotes the $d\times d$ identity matrix. Define the averages $\bar{\sigma}_{d}^{2}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d}\sum_{i=1}^{d}\sigma_{i}^{2},\,\bar{m}_{d}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d}\sum_{i=1}^{d}m_{i}.$ Denote by $\bar{m}\mathrel{\mathop{\ordinarycolon}}=\lim_{d\to\infty}\bar{m}_{d}$, assuming the limit exists in $(0,\infty)$. Then the unit-lag covariance matrix satisfies
@@
By Corollary 1.2 of Roberts et al. [13] and applying the change of variable $\boldsymbol{Y}=\boldsymbol{X}+\boldsymbol{Z}$
@@
Then fix $\boldsymbol{X}$ and set $f(\boldsymbol{X,Z})\mathrel{\mathop{\ordinarycolon}}=\alpha(\boldsymbol{X},\boldsymbol{X}+\boldsymbol{Z})$. We look at $\mathbb{E}[\alpha(\boldsymbol{X},\boldsymbol{X}+\boldsymbol{Z})\boldsymbol{Z}\boldsymbol{Z}^{\top}]$, for $\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0},\ell^{2}\boldsymbol{I}_{d}/d)$, Applying Stein’s Lemma [9] and the assumption $\mathbb{E}_{p_{i}}[|(\log p_{i})^{\prime\prime}(X_{i})|]<\infty$ gives
@@
$\tfrac{1}{2}\Big(\,\mathbb{E}[\alpha(\boldsymbol{X},\boldsymbol{X}+\boldsymbol{Z})\,\boldsymbol{Z}\boldsymbol{Z}^{\top}]-\tfrac{\ell^{2}}{d}\,\mathbb{E}[\alpha(\boldsymbol{X},\boldsymbol{X}+\boldsymbol{Z})]\boldsymbol{I}_{d}\Big)$
it follows that $\boldsymbol{R}_{d}$ is real symmetric, hence Hermitian as well. Therefore both the main diagonal term and the perturbation matrix $\boldsymbol{R}_{d}$ are Hermitian. Then by Weyl’s inequality [7], for Hermitian matrices $\boldsymbol{A}$ and $\boldsymbol{E}$,
@@
$\lambda_{\min}\!\big(\mathrm{Cov}(\boldsymbol{X}_{t},\boldsymbol{X}_{t+1})\big)\;\geq\;\min_{i}\!\Big\{\sigma_{i}^{2}-\tfrac{\ell^{2}}{d}\cdot\Phi(-\tfrac{1}{2}\ell\sqrt{\bar{m}})\Big\}-\|\boldsymbol{R}_{d}\|_{\mathrm{op}}$
Since $\min\limits_{i}\sigma_{i}^{2}>0$, the subtracted terms are of order $O(d^{-1})$, hence strictly smaller than $\min\limits_{i}\sigma_{i}^{2}$ for sufficiently large $d$. This shows that the covariance matrix is positive definite in high dimension. ∎
Theorem 4.3 shows that in the high-dimensional regime, for a product-form target with non-identically distributed components, the diagonal entries of the unit-lag covariance admit the expansion as equation (18). Roberts et al. define the following